# TestART: Improving LLM-based Unit Testing via Co-evolution of Automated Generation and Repair Iteration

### v5

@@ 4.3 Evaluation Metrics
4.4 Experimental Setup
In the experiment, TestART generates unit tests for each focal method through up to four iterations.TestART selects the best test case by prioritizing execution success, maximum coverage, and minimal test count. To compare more fairly with the baseline, TestART and ChatUniTest utilize GPT-3.5-turbo-0125 as the basic model by calling the API, which offers a 16k context length, with the temperature setting adjusted to 0.5. During the operation of ChatUniTest, we set its maxPromptTokens parameter to 16,385 and generate unit test cases through the default maxRounds of five iterations per attempt. When we conduct the baseline experiment using ChatGPT-3.5 and ChatGPT-4.0, we obtain test cases using the initial generation of TestART. We train the model of A3Test based on the Methods2Test dataset [31] and pre-trained model [32], setting the learning rate of 1e-5 for 110 epochs. We configure EvoSuite with 3 CPU cores, allocating 2000MB of memory to each core, and set the search time up to 10 minutes. During the testing process, Java 1.8 is used as the compiler and runtime environment, JUnit 4 is employed as the unit testing framework and OpenClover is used to calculate coverage rates. We calculate the mutation coverage based on the state-of-the-art mutation testing system PITest [^1] https://github.com/hcoles/pitest.git, and the mutators are set by default (11 kinds of mutators). We repeat each experiment three times and take the average as the final result.
5 Results and Analysis
@@
5.1 Answer to RQ1
TABLE II: The correctness performance of TestART compared with different baselines on Defects4J
@@ 5.1 Answer to RQ1
Figure 6: Repair rates of TestART on Defects4J
@@
Results. As presented in Table II, we first demonstrate the results of our TestART on five sub-projects respectively. Then, we compare the average results of TestART and four baseline methods on the complete dataset. The data in the table clearly shows that TestART achieves the best results with a compilation error rate of 12.43%, a runtime error rate of 8.50%, and a final pass rate of 78.55%. Compared with the second place in each indicator, TestART reduces the compilation error rate by 7.0% compared to ChatGPT-4.0 and was 8.0% lower than ChatUniTest in the running error rate. The total pass rate is 63.29% higher than A3Test, 28.64% higher than ChatGPT-3.5, 18.80% higher than ChatGPT-4.0, and 18.50% higher than ChatUniTest. We also show the repair rates and the corresponding error types in Fig. 6. The three most common compilation errors (“cannot find symbol”, “Method call error” and “Access denied”) can be repaired with about 50% success. The overall repair rate of runtime errors is as high as 75%.
@@
TABLE III: Total coverage comparison on Defects4J
@@
TABLE IV: The ablation study results of TestART on Defects4J
@@
Figure 7: Coverage rate of correct test and test case count for TestART and EvoSuite on Defects4J dataset.
5.2 Answer to RQ2
@@
Results. Table III and Fig. 7 present the sufficiency of TestART compared to the baselines. As shown in Table III, TestART achieves the highest TBC and TLC, which are 69.40% and 68.17%, respectively. Compared to the second-ranked ChatGPT-4.0, TestART improves TBC by 17.54% and TLC by 17.29%. As the same method experimented with based on the ChatGPT-3.5, ChatUniTest is not superior to ChatGPT-4.0 but exceeds the coverage of ChatGPT-3.5. Meanwhile, TestART achieves average coverage rates of 17% higher than ChatGPT-4.0. Fig. 7 shows BCCT, LCCT and the generated test case count of TestART and EvoSuite for five sub-projects. TestART achieves 88.54% branch coverage and 90.96% line coverage rates, exceeding EvoSuite by 2.5% and 3.4%, respectively. It is noteworthy that the test cases generated by TestART are about half the number produced by EvoSuite (12,359 compared to 23,570), greatly reducing the execution costs.
Analysis. Through Table III, we can see that TestART achieves the unquestionably highest coverage results compared to LLM-based methods on the total dataset. This improvement mainly benefited from the coverage-guided testing feedback and positive prompt injection. Under the coverage information guidance, incremental iteration of test cases can extensively cover areas that were missed by the original test cases. In addition, due to the inability to count the number of focal methods correctly tested by EvoSuite, we compare the coverage of the two methods on the passed focal methods. That is, the denominator of the coverage calculation is the total number of branches/lines of the focal methods correctly tested by TestART. From Fig. 7, we observe that TestART outperforms EvoSuite in three projects, while in the remaining two, it trailed by an average coverage rate of only 5%. This performance is notable given that EvoSuite, as an SBST method, seeks to cover source code by generating a large number of test cases. In contrast, TestART utilizes the powerful generative capabilities of LLMs to produce high-quality test cases in just half the quantity yet manages to achieve comparable quality to EvoSuite.
@@ 5.3 Answer to RQ3
Results. Table IV shows the ablation study results of TestART, which demonstrates that the complete TestART achieves the best performance. We present the results for four different mutations of TestART: Only ChatGPT-3.5, adding repair (+ Repair), adding repair and iteration includes positive prompt injection (+ Repair + Iteration), and adding repair, iteration, and coverage-guided testing feedback (Complete TestART). From Table IV, the repair strategy plays a vital role, increasing the Pass by 28.64%, TBC by 19.03% and TLC by 19.66%. The iteration and coverage-guided testing feedback both contribute to the improvement of the coverage rate by about 3%.
Analysis. The results from Table IV. show that different parts of TestART contribute to improving the quality of test cases. The most apparent core module for improvement is the repair strategy, which not only enhances the pass rate but also improves the coverage rate. This validates that our fixed templates do not improve pass rate at the expense of coverage. As mentioned in the motivation section, LLMs can generate high-quality test cases, but low pass rates limit the overall coverage rate. TestART solves the problem effectively through the empirical repair templates. In addition, positive prompt injection as a main part of iteration ensures that LLMs converge incrementally during loops. It is essential to suppress hallucination for LLMs; otherwise, the generated test cases will lead to a decrease in the pass rate.
@@
5.4 Answer to RQ4
Design. We set RQ4 primarily to verify TestART’s performance on unlearned new data to avoid data leakage. Since HITS [25] only experiments with certain complex focal functions within the dataset and does not validate all public focal methods of the entire project, we do not use this method as a baseline. We don’t report EvoSuite’s results on the BPG project because BPG is not executable on Java 1.8. We present the results of TestART along with two baseline methods shown in Table V. The experiment results primarily include evaluation across three dimensions: correctness (CE, RE and Pass), coverage (TBC, TLC, BCCT and LCCT) and test case count.
@@ 5.4 Answer to RQ4
Analysis. Table V clearly demonstrates that TestART maintains a high level of accuracy improvement even when facing untrained new data. The 20% increase in Pass indicates that our proposed repair strategy is equally effective on new data. We also analyze the repair results, finding that TestART successfully repaired 41.08% of compilation errors and 71.70% of runtime errors. This aligns with the repair effectiveness observed on the Defect4J dataset. In addition, The nearly 20% increase in both TBC and TLC further demonstrates that TestART achieves high testing effectiveness without significantly increasing the number of test cases. We also find that the correctness and coverage results on Defects4J and HITS dataset are almost indistinguishable, which suggests that the issue of data leakage does not significantly affect our method.
@@
TABLE V: The experimental results of TestART and two baseline methods on HITS dataset.
[…]
RUL EvoSuite - - 57.15% 27.20% 38.64% 67.81% 84.63% 421 37.40% 65.28%
ChatGPT-3.5 47.88% 20.00% 32.12% 29.94% 30.98% 93.23% 96.45% 249 32.36% 71.76%
TestART 30.30% 8.48% 59.39% 58.59% 58.14% 98.63% 98.76% 300 49.60% 75.40%
WIN EvoSuite - - 12.50% 17.39% 11.33% 73.61% 74.12% 8 11.47% 65.79%
ChatGPT-3.5 50.64% 20.51% 27.56% 26.60% 26.42% 96.51% 95.57% 200 38.07% 74.77%
TestART 30.77% 7.05% 60.26% 59.46% 58.54% 98.68% 96.19% 234 59.17% 85.43%
BPG EvoSuite - - - - - - - - - -
ChatGPT-3.5 25.89% 10.11% 63.48% 62.78% 63.05% 98.72% 98.27% 655 38.27% 87.35%
TestART 18.62% 2.84% 77.48% 75.85% 75.89% 98.74% 98.83% 7109 49.29% 88.68%
Total EvoSuite - - 35.46% 22.43% 25.37% 70.63% 79.52% - 27.90% 65.35%
ChatGPT-3.5 34.35% 13.79% 51.30% 50.28% 50.61% 97.43% 98.09% 1104 34.45% 56.94%
TestART 22.94% 4.63% 71.30% 69.74% 69.52% 98.17% 98.24% 1253 53.11% 79.20%
6 Discussion
  • Potential of error detection capability. Error detection is one of the vital abilities of unit test generation. Therefore, we adopt mutation testing to evaluate the mutation-killing ability of TestART. As shown in the last two columns of Table V, TestART achieves a total of 53.11% mutation coverage and 79.20% test strength. TestART outperforms EvoSuite by 20% in the error detection scene. This result demonstrates that TestART can effectively identify mutation errors.
@@ 6 Discussion
Figure 8: The experimental results of TestART on Internal dataset
7 Threats to Validity
@@ 8 Related Work
8.1 Automated Unit Test Generation
@@
Automated unit test generation significantly improves test efficiency. SBST [33] uses metaheuristic search techniques (such as genetic algorithms) to automate or partially automate various testing tasks. The core of this approach lies in defining a fitness function for the specific testing problem, which guides the search algorithm to find effective solutions within a potentially infinite search space. AthenaTest [1] uses a sequence-to-sequence transformer model to generate realistic, accurate and human-readable unit test cases. A3Test [5] further enhances the passing rate of AthenaTest by using a PLBART model verifying naming consistency and ensuring that test signatures match assertion knowledge. However, they suffer from low pass rates and depend excessively on fine-tuned datasets.
@@
In recent years, LLMs have performed excellently on unit test generation. In the literature [34, 35, 36, 2, 37, 38], Researchers empirically investigate the characteristics of unit test cases generated by large models and compare them with SBST tools, exploring the potential for combining the two approaches. Most methods are based on fine-tuning LLMs [39, 40, 41, 42, 43, 44] based on the code LLMs [45, 46] or are designed through prompt engineering [25, 47, 48, 49, 20, 50, 51, 52, 8, 24, 6]. As a representative work in the field, TESTPILOT [24] automatically generates unit test cases for JavaScript programs without additional training or few-shot learning. The CODAMOSA [44] proposed by Microsoft pioneers the combination of LLMs with SBST methods. ChatUnitTest [6] develop the Generation-Validation-Repair framework to create an adaptive focal context that is integrated into prompts and then submitted to ChatGPT. ChatTester [8] first understands the purpose of the focal method and then creates a unit test for it based on the help of the iterative step of intention generation. HITS [25] proposes decomposing the method-to-test into slices and generating unit tests slice by slice, applying the “divide-and-conquer” algorithm.
@@
In addition, researchers focus on the understandability [53] by contextualizing test data, enhancing identifier names, and adding descriptive comments or measuring the readability [54] of LLM-generated unit test cases. Shin et al. [55] investigate the efficacy of RAG-based LLMs in test generation. In addition to mainstream research on Java unit testing, there are also approaches for other languages. CLAP [56] is a Chat-like execution-based assert prediction for generating meaningful assert statements for Python projects. Karanjai et al. [57] explore the capabilities of generative models in crafting unit testing cases for parallel and high-performance software with C++ parallel programs. To better evaluate the unit test generation methods, scholars have proposed benchmarks [58, 59] to measure the capability of unit test generation. TestSpark [60] is a plugin for IntelliJ IDEA that enables users to generate unit tests with only a few clicks directly within their Integrated Development Environment (IDE). Among this, some researchers have also paid attention to test case repair [61, 41, 24].
@@
8.2 Automated Program Repair
In the literature [62, 63, 64], APR techniques are mainly categorized into four groups, i.e., heuristic-based [65], constraint-based [66], template-based [26], and learning-based [67] repair techniques. Our work is related to template-based APR, which is discussed below.
Template-based APR attempts to directly transform the buggy code into the correct one based on pre-defined fix patterns and represents state-of-the-art [68, 26, 69, 27]. As the flagship work in this field, TBar [26] systematically summarizes existing fix patterns and applies these patterns to patch generation. Besides, FixMiner [69] leverages a three-fold clustering strategy to extract fix patterns based on AST represent, and AVATAR [28] exploits fix pattern of static analysis tools to generate patches. Recently, inspired by the potential of combing fix patterns and LLMs, GAMMA [27] frames APR as a fill-in-the-blank task by querying LLMs to directly recover the correct code for masked cod with the code context.
Unlike template-based APR techniques, which usually focus on semantic bugs from production code, TestART attempts to repair compilation and runtime errors from test cases automatically generated by LLMs. Therefore, we design the fixed repair templates specifically in TestART for the LLM-generated unit test cases.
9 Conclusion

### v6

@@ 4.3. Evaluation Metrics
  12. (12)
    Assertion Count (AC) represents the total number of assertions contained in the generated test cases.
4.4. Experimental Setup
In the experiment, TestART generates unit tests for each focal method through up to four iterations.TestART selects the best test case by prioritizing execution success, maximum coverage, and minimal test count. To compare more fairly with the baseline, TestART and ChatUniTest utilize GPT-3.5-turbo-0125 as the basic model by calling the API, which offers a 16k context length, with the temperature setting adjusted to 0.5. During the operation of ChatUniTest, we set its maxPromptTokens parameter to 16,385 and generate unit test cases through the default maxRounds of five iterations per attempt. When we conduct the baseline experiment using ChatGPT-3.5 and ChatGPT-4.0, we obtain test cases using the initial generation of TestART. We train the model of A3Test based on the Methods2Test dataset (Tufano et al., 2020b) and pre-trained model (Ahmad et al., 2021), setting the learning rate of 1e-5 for 110 epochs. We configure EvoSuite with 3 CPU cores, allocating 2000MB of memory to each core, and set the search time up to 10 minutes. During the testing process, Java 1.8 is used as the compiler and runtime environment, JUnit 4 is employed as the unit testing framework and OpenClover is used to calculate coverage rates. We calculate the mutation coverage based on the state-of-the-art mutation testing system PITest [^1] https://github.com/hcoles/pitest.git, and the mutators are set by default (11 kinds of mutators). We repeat each experiment three times and take the average as the final result.
5. Results and Analysis
@@
5.1. Answer to RQ1 (Correctness)
Table 2. The correctness performance of TestART compared with different baselines on Defects4J
@@ 5.1. Answer to RQ1 (Correctness)
Figure 6. Repair rates of TestART on Defects4J
@@
Results. As presented in Table 2, we first demonstrate the results of our TestART on five sub-projects respectively. Then, we compare the average results of TestART and four baseline methods on the complete dataset. The data in the table clearly shows that TestART achieves the best results with a compilation error rate of 12.43%, a runtime error rate of 8.50%, and a final pass rate of 78.55%. Compared with the second place in each indicator, TestART reduces the compilation error rate by 7.0% compared to ChatGPT-4.0 and was 8.0% lower than ChatUniTest in the running error rate. The total pass rate is 63.29% higher than A3Test, 28.64% higher than ChatGPT-3.5, 18.80% higher than ChatGPT-4.0, and 18.50% higher than ChatUniTest. We also show the repair rates and the corresponding error types in Fig. 6. The three most common compilation errors (“cannot find symbol”, “Method call error” and “Access denied”) can be repaired with about 50% success. The overall repair rate of runtime errors is as high as 75%.
@@
Table 3. Total coverage comparison on Defects4J
@@
Table 4. The ablation study results of TestART on Defects4J
@@
Figure 7. Coverage rate of correct test and test case count for TestART and EvoSuite on Defects4J dataset.
5.2. Answer to RQ2 (Sufficiency)
@@
Results. Table 3 and Fig. 7 present the sufficiency of TestART compared to the baselines. As shown in Table 3, TestART achieves the highest TBC and TLC, which are 69.40% and 68.17%, respectively. Compared to the second-ranked ChatGPT-4.0, TestART improves TBC by 17.54% and TLC by 17.29%. As the same method experimented with based on the ChatGPT-3.5, ChatUniTest is not superior to ChatGPT-4.0 but exceeds the coverage of ChatGPT-3.5. Meanwhile, TestART achieves average coverage rates of 17% higher than ChatGPT-4.0. Fig. 7 shows BCCT, LCCT and the generated test case count of TestART and EvoSuite for five sub-projects. TestART achieves 88.54% branch coverage and 90.96% line coverage rates, exceeding EvoSuite by 2.5% and 3.4%, respectively. It is noteworthy that the test cases generated by TestART are about half the number produced by EvoSuite (12,359 compared to 23,570), greatly reducing the execution costs.
Analysis. Through Table 3, we can see that TestART achieves the unquestionably highest coverage results compared to LLM-based methods on the total dataset. This improvement mainly benefited from the coverage-guided testing feedback and positive prompt injection. Under the coverage information guidance, incremental iteration of test cases can extensively cover areas that were missed by the original test cases. In addition, due to the inability to count the number of focal methods correctly tested by EvoSuite, we compare the coverage of the two methods on the passed focal methods. That is, the denominator of the coverage calculation is the total number of branches/lines of the focal methods correctly tested by TestART. From Fig. 7, we observe that TestART outperforms EvoSuite in three projects, while in the remaining two, it trailed by an average coverage rate of only 5%. This performance is notable given that EvoSuite, as an SBST method, seeks to cover source code by generating a large number of test cases. In contrast, TestART utilizes the powerful generative capabilities of LLMs to produce high-quality test cases in just half the quantity yet manages to achieve comparable quality to EvoSuite.
@@ 5.3. Answer to RQ3 (Ablation)
Results. Table 4 shows the ablation study results of TestART, which demonstrates that the complete TestART achieves the best performance. We present the results for four different mutations of TestART: Only ChatGPT-3.5, adding repair (+ Repair), adding repair and iteration includes positive prompt injection (+ Repair + Iteration), and adding repair, iteration, and coverage-guided testing feedback (Complete TestART). From Table 4, the repair strategy plays a vital role, increasing the Pass by 28.64%, TBC by 19.03% and TLC by 19.66%. The iteration and coverage-guided testing feedback both contribute to the improvement of the coverage rate by about 3%.
Analysis. The results from Table 4. show that different parts of TestART contribute to improving the quality of test cases. The most apparent core module for improvement is the repair strategy, which not only enhances the pass rate but also improves the coverage rate. This validates that our fixed templates do not improve pass rate at the expense of coverage. As mentioned in the motivation section, LLMs can generate high-quality test cases, but low pass rates limit the overall coverage rate. TestART solves the problem effectively through the empirical repair templates. In addition, positive prompt injection as a main part of iteration ensures that LLMs converge incrementally during loops. It is essential to suppress hallucination for LLMs; otherwise, the generated test cases will lead to a decrease in the pass rate.
@@
5.4. Answer to RQ4 (Generalization)
Table 5. The experimental results of TestART and two baseline methods on HITS dataset.
[…]
RUL EvoSuite - - 57.15% 27.20% 38.64% 67.81% 84.63% 421 37.40% 65.28% -
ChatGPT-3.5 47.88% 20.00% 32.12% 29.94% 30.98% 93.23% 96.45% 249 32.36% 71.76% 181
TestART 30.30% 8.48% 59.39% 58.59% 58.14% 98.63% 98.76% 300 49.60% 75.40% 320
WIN EvoSuite - - 12.50% 17.39% 11.33% 73.61% 74.12% 8 11.47% 65.79% -
ChatGPT-3.5 50.64% 20.51% 27.56% 26.60% 26.42% 96.51% 95.57% 200 38.07% 74.77% 92
TestART 30.77% 7.05% 60.26% 59.46% 58.54% 98.68% 96.19% 234 59.17% 85.43% 260
BPG EvoSuite - - - - - - - - - - -
ChatGPT-3.5 25.89% 10.11% 63.48% 62.78% 63.05% 98.72% 98.27% 655 38.27% 87.35% 858
TestART 18.62% 2.84% 77.48% 75.85% 75.89% 98.74% 98.83% 7109 49.29% 88.68% 1120
Total EvoSuite - - 35.46% 22.43% 25.37% 70.63% 79.52% - 27.90% 65.35% -
ChatGPT-3.5 34.35% 13.79% 51.30% 50.28% 50.61% 97.43% 98.09% 1104 34.45% 56.94% 1131
TestART 22.94% 4.63% 71.30% 69.74% 69.52% 98.17% 98.24% 1253 53.11% 79.20% 1700
Design. We set RQ4 primarily to verify TestART’s performance on unlearned new data to avoid data leakage. Since HITS (Wang et al., 2024b) only experiments with certain complex focal functions within the dataset and does not validate all public focal methods of the entire project, we do not use this method as a baseline. We don’t report EvoSuite’s results on the BPG project because BPG is not executable on Java 1.8. We present the results of TestART along with two baseline methods shown in Table 5. The experiment results primarily include evaluation across three dimensions: correctness (CE, RE and Pass), coverage (TBC, TLC, BCCT and LCCT) and test case count.
@@ 5.4. Answer to RQ4 (Generalization)
Analysis. Table 5 clearly demonstrates that TestART maintains a high level of accuracy improvement even when facing untrained new data. The 20% increase in Pass indicates that our proposed repair strategy is equally effective on new data. We also analyze the repair results, finding that TestART successfully repaired 41.08% of compilation errors and 71.70% of runtime errors. This aligns with the repair effectiveness observed on the Defect4J dataset. In addition, The nearly 20% increase in both TBC and TLC further demonstrates that TestART achieves high testing effectiveness without significantly increasing the number of test cases. We also find that the correctness and coverage results on Defects4J and HITS dataset are almost indistinguishable, which suggests that the issue of data leakage does not significantly affect our method.
@@
(a) Time cost of TestART for single focal method.
(b) Money cost of TestART for 100 generations of each project.
Figure 8. Cost of TestART.
5.5. Answer to RQ5 (Cost)
Design. In software engineering, both time cost and monetary cost are crucial metrics for evaluating unit test generation tools. In terms of time cost, if the generation process takes too long, it will negatively impact development efficiency and reduce the tool’s practicality. As for monetary cost, TestART utilizes ChatGPT by calling APIs, and each invocation incurs a certain fee. Whether this cost is reasonable is also a significant factor affecting the practicality of the unit test generation tool. This experiment evaluates the temporal efficiency of TestART by measuring the time required to generate unit test cases for individual focal methods. Additionally, it quantifies the monetary expenditure across three key phases—initial generation, repair, and iteration—for each project within the Defects4J benchmark.
Results and Analysis. Fig. 8a illustrates the time consumption of TestART for unit test generation tasks. As shown in the figure, 68.58% of the unit test generation tasks are completed within one minute, while 90.59% are finished within two minutes. Although a small portion of tasks requiring repair and iteration may consume more time, 99% of the tasks can be completed within five minutes. The maximum search time we set when generating test cases using EvoSuite is 10 minutes, which means TestART is more efficient than EvoSuite.
Fig. 8b presents the average cost of TestART for every 100 generations across different projects. The experiment calculates the cost based on the pricing of GPT-3.5-turbo-0125: 0.5\$/Mtoken for prompt and 1.5\$/Mtoken for completion. As shown in the figure, the total cost varies among different projects. Gson has the lowest unit test generation cost, averaging approximately 0.2\$ for 100 times unit test generation, while Lang requires about 0.45\$. Furthermore, the initial generation phase accounts for the largest proportion of the total cost, constituting approximately 50%. The combined costs of code repair and iterative feedback make up the remaining 50%. In some projects, such as Lang, the repair cost is higher than the iteration cost. This is because most focal methods in Lang are relatively simple, and once the test cases pass execution, they can achieve full branch and statement coverage without requiring further iterations. In contrast, for the Chart project, the iteration cost exceeds the code repair cost. This is attributed to the higher proportion of complex and large focal methods in this project, which often require multiple iterations to achieve high coverage, thereby increasing the cost of iterative feedback.
Answer to RQ5 TestART demonstrates acceptable and controllable time and money cost, with 90.59% of a single generation task completed within two minutes and the money cost for 100 generation tasks ranging from 0.3\$-0.45\$
Figure 9. The experimental results of TestART on Internal dataset
6. Discussion
  • Potential of error detection capability. Error detection is one of the vital abilities of unit test generation. Therefore, we adopt mutation testing to evaluate the mutation-killing ability of TestART. As shown in the MC and TS columns of Table 5, TestART achieves a total of 53.11% mutation coverage (MC) and 79.20% test strength (TS). TestART outperforms EvoSuite by 20% in the error detection scene. This result demonstrates that TestART can effectively identify mutation errors.
  • Quality of assertion generation. The number of assertions in a test case is important for measuring the quality of the test case. As shown in the last column of Table 5, the total number of assertions generated by TestART exceeds that of ChatGPT-3.5 by 569. This demonstrates that TestART not only improves the code coverage of test cases but also maintains a substantial number of assertions, ensuring comprehensive test validation.
@@ 6. Discussion
7. Threats to Validity
@@ 8. Related Work
8.1. Automated Unit Test Generation
@@
Automated unit test generation significantly improves test efficiency. SBST (McMinn, 2004) uses metaheuristic search techniques (such as genetic algorithms) to automate or partially automate various testing tasks. The core of this approach lies in defining a fitness function for the specific testing problem, which guides the search algorithm to find effective solutions within a potentially infinite search space. AthenaTest (Tufano et al., 2020a) uses a sequence-to-sequence transformer model to generate realistic, accurate and human-readable unit test cases. A3Test (Alagarsamy et al., 2024a) further enhances the passing rate of AthenaTest by using a PLBART model verifying naming consistency and ensuring that test signatures match assertion knowledge. However, they suffer from low pass rates and depend excessively on fine-tuned datasets.
@@
In recent years, LLMs have performed excellently on unit test generation. In the literature (Xiao et al., 2024; Yang et al., 2024b; Schäfer et al., 2024; Tang et al., 2024; Huang et al., 2024; Siddiq et al., 2023), Researchers empirically investigate the characteristics of unit test cases generated by large models and compare them with SBST tools, exploring the potential for combining the two approaches. Most methods are based on fine-tuning LLMs (Dakhel et al., 2024; Ni et al., 2024; Yaraghi et al., 2024; Plein et al., 2024; Shin et al., 2024b; Lemieux et al., 2023) based on the code LLMs (Le et al., 2022; Wang et al., 2021) or are designed through prompt engineering (Wang et al., 2024b; Yang et al., 2024a; Pizzorno and Berger, 2024; Alagarsamy et al., 2024b; Alshahwan et al., 2024; Ryan et al., 2024; Jiang et al., 2024; Li and Doiron, 2023; Yuan et al., 2024; Schäfer et al., 2023; Chen et al., 2024). As a representative work in the field, TESTPILOT (Schäfer et al., 2023) automatically generates unit test cases for JavaScript programs without additional training or few-shot learning. The CODAMOSA (Lemieux et al., 2023) proposed by Microsoft pioneers the combination of LLMs with SBST methods. ChatUnitTest (Chen et al., 2024) develop the Generation-Validation-Repair framework to create an adaptive focal context that is integrated into prompts and then submitted to ChatGPT. ChatTester (Yuan et al., 2024) first understands the purpose of the focal method and then creates a unit test for it based on the help of the iterative step of intention generation. HITS (Wang et al., 2024b) proposes decomposing the method-to-test into slices and generating unit tests slice by slice, applying the “divide-and-conquer” algorithm. TestART demonstrates distinct advantages over ChatUnitTest (Chen et al., 2024) and ChatTester (Yuan et al., 2024) in both repair strategies and the iterative process. First, TestART’s repair mechanism employs carefully designed templates derived from automated program repair (APR) research, enabling efficient automated repair of typical compilation and runtime errors in LLM-generated test cases. In contrast, ChatTester relies solely on unstable LLM-based error correction, while ChatUnitTest employs only basic syntactic and import repair rules. Second, TestART implements a coverage-driven iterative process that incrementally generates test cases through the co-evolution of generation and repair. This differs fundamentally from ChatUnitTest’s non-iterative approach and ChatTester’s feedback mechanism neglects coverage-guided information.
@@
In addition, researchers focus on the understandability (Deljouyi et al., 2024) by contextualizing test data, enhancing identifier names, and adding descriptive comments or measuring the readability (Zhou et al., 2024) of LLM-generated unit test cases. Shin et al. (Shin et al., 2024a) investigate the efficacy of RAG-based LLMs in test generation. In addition to mainstream research on Java unit testing, there are also approaches for other languages. CLAP (Wang et al., 2024a) is a Chat-like execution-based assert prediction for generating meaningful assert statements for Python projects. Karanjai et al. (Karanjai et al., 2024) explore the capabilities of generative models in crafting unit testing cases for parallel and high-performance software with C++ parallel programs. To better evaluate the unit test generation methods, scholars have proposed benchmarks (Zhang et al., 2024; Lops et al., 2024) to measure the capability of unit test generation. TestSpark (Sapozhnikov et al., 2024) is a plugin for IntelliJ IDEA that enables users to generate unit tests with only a few clicks directly within their Integrated Development Environment (IDE). Among this, some researchers have also paid attention to test case repair (Liu et al., 2024; Yaraghi et al., 2024; Schäfer et al., 2023).
@@
8.2. Automated Program Repair
In the literature, APR techniques are mainly categorized into four groups (Zhang et al., 2023a; Gazzola et al., 2019; Monperrus, 2018), i.e., heuristic-based (Le Goues et al., 2012), constraint-based (Xiong et al., 2017), template-based (Liu et al., 2019b), and learning-based (Yuan et al., 2022) repair techniques. Our work is related to template-based APR, which is discussed below.
Template-based APR attempts to directly transform the buggy code into the correct one based on pre-defined fix patterns and represents state-of-the-art (Monperrus, 2014; Liu et al., 2019b; Koyuncu et al., 2020; Zhang et al., 2023b). As the flagship work in this field, TBar (Liu et al., 2019b) systematically summarizes existing fix patterns and applies these patterns to patch generation. Besides, FixMiner (Koyuncu et al., 2020) leverages a three-fold clustering strategy to extract fix patterns based on AST represent, and AVATAR (Liu et al., 2019a) exploits fix pattern of static analysis tools to generate patches. Recently, inspired by the potential of combing fix patterns and LLMs, GAMMA (Zhang et al., 2023b) frames APR as a fill-in-the-blank task by querying LLMs to directly recover the correct code for masked cod with the code context. According to ReAssert (Daniel et al., 2009), repairing unit tests by empirical strategies is effective and acceptable for developers. However, this strategy is not for test cases generated from LLMs.
Therefore, to increase test case passing rates, TestART designs a template-driven repair strategy that is adaptive to LLM outputs inspired by the APR community. Unlike template-based APR techniques, which usually focus on semantic bugs from production code, TestART attempts to design simple and efficient templates to repair the typical compilation and runtime errors from test cases automatically generated by LLMs. More importantly, TestART implements the co-evolution between the generation and repair process to generate high-quality test cases under several iterations.
9. Conclusion